OpenStudy (anonymous):

Suppose you place 2 black marbles, 3 yellow marbles, and 5 red marbles in a bag. Find the probability of pulling a red then a yellow? The marbles ARE replaced after drawn.

jigglypuff314 (jigglypuff314):

Hello @Boby6579 and Welcome to OpenStudy! :) Probability of a single event is the (number of chances of it happening) over (number of total chances overall) total in this case is 2+3+5 = 10 and the chance of red = 5 and the chance of yellow = 3 so the chance of red by itself is 5/10 = 1/2 and the chance of yellow by itself is 3/10 Follow so far? :)

jigglypuff314 (jigglypuff314):

then, the probabilty of both happening is the chances of them Multiplied together so 1/2 times 3/10 = answer

OpenStudy (anonymous):

thanks